I need help with this problem: A pound of tin is worth $1 more than a pound of copper. Three pounds of tin are mixed with 7 pounds of copper to make bronze that sells for $3.05 per pound. How much is a pound of tin worth?
Answer by stanbon(75887)

A pound of tin is worth $1 more than a pound of copper. Three pounds of tin are mixed with 7 pounds of copper to make bronze that sells for $3.05 per pound. How much is a pound of tin worth?
------------------------
Let copper be worth $x per pound
Then tin is worth $(x+1) per pound.
-----
Equation:
Value Equation: 3(x+1) + 7x = 3.05*10
3x+3 + 7x = 30.50
10x = 27.50
---
x = $2.75 (price of a pound of copper)
---
x+1 = $3.75 (price of a pound of tin)
======================================
Cheers,
Stan H.
============